Database last updated on Fri, 06 Feb 2026 02:10:53 +0000 by localhost

Critical warnings from b-files' comments for days between 2022069 and 2022069 (0 sec, 6 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
td:Date changed from 10/03/22 to 11/03/2211B06922.204TORON#20402:19:501-1
rl:2022 070 03 37 01 Motor #2: appears stuck with a reference sensor activated.11B06922.204TORON#20405:29:531-1
rl:2022 069 13 11 34 Bad command string. Command string buffer cleared.1093B06922.204TORON#20413:36:541-1
rl:2022 069 13 11 34 Bad command string. Command string buffer cleared.1093B06922.204TORON#20413:36:541-1
hg:Lamp not on... hg test terminated F1= 1 F4= 1621B06922.204TORON#20416:00:491-1
hg:Lamp not detected f1= 1 f4= 1621B06922.204TORON#20416:00:501-1

Instrument constant changes for the week before 2022069 (2 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
204202206952913.0

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
20420220691280.0

Non-critical warnings from b-files' comments for days between 2022069 and 2022069 (0 sec, 8 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 068 02 42 39 Attempt to exceed min boundary13B06922.204TORON#20405:29:551-1
rl:2022 068 02 42 39 <M,9,-1746>13B06922.204TORON#20405:29:571-1
rl:2022 068 02 42 39 <M,9,-1746>15B06922.204TORON#20405:29:571-1
dp:Doing dispersion with internal HG.199B06922.204TORON#20410:59:471-1
hg:DEL key pressed199B06922.204TORON#20413:33:511-1
checkhg:Last HG had to adjust the micrometer by at least 98399B06922.204TORON#20417:11:151-1
uv:UV extended scan aborted13B06922.204TORON#20419:56:141-1
fr:Micrometer #1 reference found at step 8950 (+CAL)= 9863 Now set to o3 offset = 7303199B06922.204TORON#20420:59:29986310009000

Measurements and tests summary.


_jday_loc_serial_skcedhghpuv
069TORONTO#204menu,Waiting until ZA = 170 now 127.6,Waiting until ZA = 170 now 129,Waiting until ZA = 170 now 130.3,Waiting until ZA = 170 now 131.6,Waiting until ZA = 170 now 132.8,Waiting until ZA = 170 now 133.9,Waiting until ZA = 170 now 135,Waiting until ZA = 170 now 136,Waiting until ZA = 170 now 136.8,Waiting until ZA = 170 now 137.6,Waiting until ZA = 170 now 138.3,Waiting until ZA = 170 now 138.8,Waiting until ZA = 170 now 139.3,Waiting until ZA = 170 now 139.6,Waiting until ZA = 170 now 139.8,Waiting until ZA = 170 now 139.9237481
(0 sec)

Relative and absolute humidity for day 2022069 (2 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
20418.24.2